OpenStudy (anonymous):

thurman put $90 into an account that pays 4.4% interest compounded quarterly. according to the rule of 72 approximately how long will it take for his money to double

OpenStudy (ranga):

I think it should be 72/4.4 = 16 years (approx). If you want to take into account the quarterly compounding, then the interest rate per quarter is 4.4/4 = 1.1 Then the rule of 72 says the principal will double in 72/1.1 quarters or 72/1.1 * 1/4 years = 72/4.4 = 16 years (approx). Since its is an estimate, the compounding period does not enter the picture when using the rule of 72.